Please solve 6.14, problem 6.13 included for context
Answer #1

31n k-167m r)-プ gmm s 31mm 下 12 @w= (tas-5)(742)(i4s5)HEPY132) (İaS3t13 2.) 23-2 14S3 4 13-2 = 415200mm 9o 11.5x 7-62